From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Jan 11 2003 - 16:03:48 EST
------------------------------------------------------------ Date: Sat Jan 11 16:03:47 2003, Shift: 08:00-16:00 Supervisor: JIJ, with: RP Log book#13 pages 51 - 52 Shift Summary: ------------------------------------------------------------ 1. The machine performance was stable although with lots of background. Scalers 15 (H1) and 16 (H2) had initial rates of 300k and 400k per second, respectively. (Typical AuAu rates 2001 were 15k). At the end of the store rates were down to the order of 50k and 70k respectively. 2. 7324 - DVM run 7326 - 95k tr. 6, T3 and T5 off 7327 - 38k tr. 6 T3 section 1 on/off 7328 - Pedestal run 7329 - TPC pedestals 3. T3 had high current and should stay off untill next acces to resolve wire plane coupling problem. T4 works fine. T5 had large current while high background. Initially the current was ~15-25uA and finally a few uA. H2 reported with poor performance in owl shift now seems to work fine. Reason for remedy unknown. MRS trigger (trigger 3) still picks up lots of bad events seen from BB vertex distribution. In data taking it was scaled down additionally. T1 had high current and anodes were turned off by request of Ramiro. T2, TPM1 and TPM2 show current oscillations. Ramiro suspects space charge buildup from high background. ------------------------------------------------------------
This archive was generated by hypermail 2.1.5 : Sat Jan 11 2003 - 16:06:38 EST